OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7

OpenStudy (anonymous):

\[ x = 4py^2 \]

OpenStudy (anonymous):

Where \(p\) is the the \(y\) of the focus minus the \(y\) of the vertex.

OpenStudy (anonymous):

The vertex is halfway between focus and directrix.

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

so is it a sideways parabola?

OpenStudy (anonymous):

Yeah it looks like it.

OpenStudy (anonymous):

so p=3 but how do i know if the focus is (3,0) or (0,3)?

OpenStudy (anonymous):

no, \(p=7\)

OpenStudy (anonymous):

wrong problem lol. i waslooking at another one

OpenStudy (anonymous):

so this parabola has the vertex at (0,0) and then what do i do from then on?

OpenStudy (anonymous):

All you ever have to do is find \(p\).

OpenStudy (anonymous):

and figure out if it is horizontal or vertical

OpenStudy (anonymous):

You can use:\[\Large p = \frac{x_{focus}-x_{directrix}}{2} \] or\[\Large p = x_{focus}-x_{vertex} \]

OpenStudy (anonymous):

When your directrix is \(x =b\), then you have \(x=4py^2\) When your directrix is \(y=b\) then you have \(y=4px^2\).

OpenStudy (anonymous):

You can can translate it by \((x_0,y_0)\), using \((x,y)\to (x-x_0,y-y_0)\)

OpenStudy (anonymous):

@wio i got x=28y^2 but that's not one of the answer choices

OpenStudy (anonymous):

what are the answer choices

OpenStudy (anonymous):

y = one divided by twenty eightx2 x = one divided by twenty eighty2 -28y = x2 y2 = 14x

OpenStudy (anonymous):

y=(1/28)x2 x=(1/28)y2 are the first and second choirces my bad

OpenStudy (anonymous):

the 2 is squared

OpenStudy (anonymous):

My bad it was \(4px = y^2\)

OpenStudy (anonymous):

So you should get \(1/28\)

OpenStudy (anonymous):

\When your directrix is \(x =b\), then you have \(4px=y^2\) When your directrix is \(y=b\) then you have \(4py=x^2\).

OpenStudy (anonymous):

oh okay, THENK YOU SO MUCH